OpenStudy (anonymous):

help please ! -4x+-2=3x-10 i did all the work before i'm just not sure what to do after

jimthompson5910 (jim_thompson5910):

Your next step is to add 2 to both sides. What do you get?

OpenStudy (anonymous):

-8

jimthompson5910 (jim_thompson5910):

so -4x+-2=3x-10 becomes -4x=3x-8

jimthompson5910 (jim_thompson5910):

Now subtract 3x from both sides

OpenStudy (anonymous):

so it would be -x=8

OpenStudy (anonymous):

-8

jimthompson5910 (jim_thompson5910):

good, x = -8 is your final answer

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

oh wait, one sec

jimthompson5910 (jim_thompson5910):

it should be -x=-8 which turns into x = 8

OpenStudy (anonymous):

oh.. okay ! thank you

OpenStudy (anonymous):

can you help me with another one...

jimthompson5910 (jim_thompson5910):

hold on, let me check something real fast

OpenStudy (anonymous):

:p okay

OpenStudy (anonymous):

Isn't the answer actually x = 8/7?

OpenStudy (anonymous):

thats not one of my options...

jimthompson5910 (jim_thompson5910):

oh my bad, i made an error, here's the correct way to do this -4x+-2=3x-10 -4x=3x-10+2 -4x = 3x - 8 -4x-3x = - 8 -7x = -8 x = -8/(-7) x = 8/7 Not sure how I missed that, but oh well

OpenStudy (anonymous):

would you like to see the beginning?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

6x − 2(x + 1) = 3x − 10

jimthompson5910 (jim_thompson5910):

6x − 2(x + 1) = 3x − 10 6x − 2x - 2 = 3x − 10 4x - 2 = 3x − 10 4x = 3x − 10 + 2 4x = 3x − 8 4x -3x = -8 1x = -8 x = -8

jimthompson5910 (jim_thompson5910):

So the error you made was thinking that 6x - 2x was -4x it should be 6x - 2x = 4x

OpenStudy (anonymous):

oh.. im sorry :p

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

can you help me with another one

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

After a discount of 20%, the price of a pair of shoes, including tax, is $66.91. The following equation can be used to find x, the price of the pair of shoes without the discount. x - 0.2x + 0.06(x - 0.2x) = 66.91 What is the price of the pair of shoes without the discount?

jimthompson5910 (jim_thompson5910):

They want you to solve for x

jimthompson5910 (jim_thompson5910):

x - 0.2x + 0.06(x - 0.2x) = 66.91 x - 0.2x + 0.06(0.8x) = 66.91 x - 0.2x + 0.048x = 66.91 0.848x = 66.91 Can you take it from here?

OpenStudy (anonymous):

yes thank you

jimthompson5910 (jim_thompson5910):

yw